A volume of 18.0 L contains a mixture of 0.250 mol
Answer #1

For mixture, Volume = V = 18.0 L

temperature = T = 0 oC = 273 K

Pressure = P = 1 atm

R = 0.082 L atm/mol K

Applyting formula,

PV= nRT

n = PV/RT = (1)(18.0)/(0.082*273) = 0.804 mol

n = total number of moles = 0.804

n = number of moles of (N2 + O2 + He)

0.804 = 0.250 + 0.250 + number of moles of He

number of moles of He = 0.804-0.500 = 0.304 mol

Molar mass of He = 4 g/mol

mass of He = molar mass*number of moles = 4*0.304 = 1.216 g = 1.22 g

mass of He present in mixture = 1.22 g